# Cosmological simulations of the generation of cluster-scale radio emission from turbulent re-acceleration

### v1

@@ Abstract
Aims. We investigate whether the formation mechanisms of radio halos and mega radio halos differ, or whether they can be produced by different modalities of the same (re)acceleration mechanism. Here we present results of a cosmological simulation of a disturbed galaxy cluster, with the aim to study the origin of mega radio halos.
Methods. We analysed the evolution of cosmic-ray electrons, subject to gains and losses using a Fokker-Planck solver. In particular, we included the effects of Adiabatic Stochastic Acceleration (ASA) which is caused by the stochastic interaction of cosmic rays with diffusing magnetic field lines in super-Alfvenic turbulence. Moreover, we included shock acceleration and the seeding of cosmic-ray electrons by galaxies.
@@ 1 Introduction
For at least three decades it has been known that galaxy clusters host diffuse radio sources that are not associated with individual cluster galaxies (e.g. van Weeren et al., 2019, for a review). Such extended radio sources are associated with a perturbed dynamical state of the host cluster, and probe the existence of cosmic rays and magnetic fields in the intracluster medium (e.g. Brunetti & Jones, 2014, for a review). The unprecedented sensitivity to large-scale emission provided by the Low Frequency Array (LOFAR, van Haarlem et al. 2013) has led to the discovery of very extended radio emission in clusters and between pairs of clusters (e.g. Govoni et al., 2019; Botteon et al., 2020; Osinga et al., 2021; Hoeft et al., 2021; Bonafede et al., 2021). Botteon et al. (2022) reported the LOFAR detection of diffuse radio emission in Abell 2255, covering almost the entire cluster volume, and containing classical radio halos and radio relics. At the same time, Cuciti et al. (2022) discovered a new class of diffuse radio sources beyond the volume of classical radio halos, dubbed mega radio halos (hereafter MegaRHs, in contrast to the ClassicalRHs introduced before). In four clusters, the ClassicalRH was found to be embedded in a much larger source that extends to scales of 2-3 Mpc, filling the volume of the cluster, at least up to $R_{500}$. The radial profiles of the radio emission in these newly discovered MegaRHs show two distinct components: a brighter core region that corresponds to the ClassicalRH and, separated by a clear discontinuity, a more extended and shallower low-surface brightness component. Based on the small number of known MegaRHs, the transition from the ClassicalRHs to MegaRHs occurs at a radius which is about a half of $R_{\rm 500}$ from the cluster centre. The average emissivity of the MegaRHs is $\sim 20-25$ times lower than the emissivity of ClassicalRHs. For two MegaRHs, the combination of 50 and 140 MHz LOFAR observations yielded a relatively steep spectrum with a spectral index $\alpha\sim-1.6$.
@@
MegaRHs show promise to reveal the physical properties of the outskirts of galaxy clusters, in a $\sim 30$ times larger volume than what has been possible to date. They may also help understand particle acceleration and magnetic field amplification in very weakly collisional plasmas (e.g. Brunetti & Lazarian, 2011; Brunetti & Jones, 2014; Rincon et al., 2016; Sironi et al., 2023; Zhou et al., 2023; St-Onge & Kunz, 2018; Kunz et al., 2022).
@@
Cluster-wide radio emission morphologies are compatible with a scenario where relativistic electrons are re-accelerated by second-order Fermi mechanisms in the turbulent intracluster medium, ICM (e.g. Brunetti et al., 2001; Fujita et al., 2003; Cassano & Brunetti, 2005; Brunetti & Lazarian, 2016). In MegaRH, in particular, their steep radio spectrum makes this explanation by far the most plausible (Cuciti et al., 2022; Beduzzi et al., 2023; Nishiwaki et al., 2024).
@@
The theoretical challenge here is to maintain the relativistic electrons at energies that keep their synchrotron emission detectable at radio frequencies. The discontinuity in the surface brightness profile suggests a change either in the macrophysical (e.g. acceleration mechanisms) or in the microphysical properties of the plasma (e.g. acceleration efficiency, mean free path) when moving from the ClassicalRH to the MegaRH region. The scenario that we explore here is that, both, MegaRH and ClassicalRH are powered by turbulence (e.g. Brunetti et al., 2007; Brunetti & Jones, 2014), but their observable difference follow the significantly different properties of turbulence in the two regions. Previous studies, using cosmological simulations to interpret real observations (Botteon et al., 2022; Nishiwaki et al., 2024) suggested that radio emission in peripheral cluster regions require several percent of dissipation of kinetic turbulent energy into the amplification of magnetic fields and into the acceleration of relativistic electrons. This can happen if the turbulence in the MegaRH region has different properties than in the ClassicalRH region, and this work is also devoted to explore these possible differences.
@@
In a first attempt (Beduzzi et al., 2023, hereafter Paper I) we investigated the lifecycle of relativistic electrons in the outskirts of galaxy clusters under realistic conditions. We calculated under which circumstances the observed volume of MegaRHs could be filled with radio-emitting electrons. However, in this first work we did not model the full ageing and re-acceleration process of cosmic-ray electrons (CRe) mixed with the ICM, nor did we simulate their synchrotron emission in any detail. Here we extend our previous work, this in order to quantitatively test the theoretical formation scenarios of MegaRH.
@@
A few other papers recently explored the formation of very extended radio emission in simulated clusters of galaxies, albeit with a more simplified approach than the one we are going to use in this work. By relying on the ”classical” Transit Time Damping (TTD) acceleration (in which the re-acceleration is powered by magneto-sonic waves excited by the compressive part of the turbulent kinetic power e.g. Cassano & Brunetti 2005; Brunetti et al. 2007; Donnert et al. 2013) Böss et al. (2023a) simulated the emergence of diffuse radio emission in in a suite of constrained simulations of the Local Universe to study the production of radio emission from turbulent re-acceleration in massive clusters, such as Coma. On the other hand Nishiwaki et al. (2024), explored the adiabatic-stochastic-(re)acceleration scenario (Brunetti & Lazarian, 2011; Brunetti & Vazza, 2020) and used a Fokker-Planck method to evolve the distribution of relativistic electrons in a perturbed cluster from the same suite of MHD cosmological simulations used in our Paper I. Limited to the analysis of a single cluster snapshot from the simulation, they found that if a conservative fraction ($\sim 1\%$) of solenoidal turbulent energy is dissipated into magnetic energy and electron re-acceleration, a large fraction of peripheral cluster volume can produce detectable radio emission.
@@
In this work, we apply a full Fokker-Planck solver over a population of Lagrangian tracers. This will be done for about a hundred root-grid time steps (approximately yielding a $\Delta t=100\leavevmode\nobreak\ \rm Myr$ time step) of the simulation in order to test whether the complex cycles of (re)acceleration experienced by relativistic electrons can reproduce observational properties of MegaRH.
@@ 2.1 Cosmological simulation: ENZO
Continuing on from Paper I, we analysed the same cosmological simulation of a cluster of galaxies, produced with ENZO code, for which our previous work predicted significant chances of re-acceleration of relativistic electrons, in both the Classical and Mega RH regions. The simulation includes the self-gravity for ordinary and dark matter, radiative equilibrium cooling, but no other galaxy formation-related physics.
No feedback processes were needed as the resolution is not high enough to produce strong cooling flow, and the cluster is undergoing frequent merging to keep it from cooling castrophically. Magnetic fields are evolved using ideal magnetohydrodynamics (MHD) using the hyperbolic cleaning method (Bryan et al., 2014). At the beginning of the simulation, at a redshift of $z=40$, we injected a uniform magnetic field of $\rm B_{0}=0.4$ nG in all directions. The most massive cluster at $z=0$ has a mass of $\rm M_{100}=3.8\times 10^{14}\ M_{\odot}$ and a virial radius of $\rm R_{100}=1.5$ Mpc. The simulation has been performed on a computational grid of $128^{3}$ cells, namely $(100\rm\leavevmode\nobreak\ Mpc/h^{-1})^{3}$, with dark matter particles ($\rm m_{DM0}=3\times 10^{10}M_{\odot}$ of mass resolution). A zoomed-in region where the cluster forms (with size $(16\,{\rm Mpc})^{3}$) was further refined with four additional levels of ×$\times 2$ refinement in spatial resolution each (and ×$\times 2^{3}$ refinement in mass resolution of the dark matter component, at each refinement step) with nested regions of decreasing size. The entire cluster region is thus resolved with a maximum dark matter mass resolution of $m_{\rm DM}=7.3\times 10^{6}M_{\odot}$ (i.e. $m_{\rm DM0}/8^{4}$), and with a uniform spatial resolution of $70$ kpc/cell. On top of this, we also allowed for two extra levels of adaptive mesh refinement within the innermost region by refining on local gas or dark matter overdensities, up to a final maximum resolution of $\approx 12\rm\leavevmode\nobreak\ kpc/h=18$ kpc (comoving).
@@
The assumed cosmology is a flat $\Lambda$CDM model with $H_{0}=67.8\,\rm km/s/Mpc$, $\Omega_{M}=0.308$, $\Omega_{\Lambda}=0.692$, $\Omega_{b}=0.0478$ and $\sigma_{8}=0.811$.
@@
Figure 2: Evolving momentum spectra of all populations of relativistic electrons, for three physical models (C=only cooling; CT=cooling and turbulent re-acceleration; CST=cooling, turbulent-reacceleration and shock acceleration). The panels show the same times as in Fig. 1 (with the addition of the initial time step after the first injection of electrons in tracers): 3.38, 5.86, 8.58, 9.66, 10.61, 11.36, 12.14, and 12.64 Gyr.
@@ 2.2 Simulations of cosmic-ray electron spectra: CRATER & ROGER
We modelled the spatial evolution of relativistic electrons in post-processing by sampling of a distribution of discrete Lagrangian particles, created by sampling a fixed gas mass resolution (approximately of $5\cdot 10^{8}M_{\odot}$) in the grid of density values produced by the ENZO simulation. Once injected, the trajectories of tracers are evolved forward in time, with a simple time integrator ${r}(t+dt)=\@vec{r}(t)+\@vec{v_{\rm gas}}dt$, after interpolating the 3d-velocity field with a cloud-in-cell spatial interpolation method, as explained in Wittor et al. (2017). To each Lagrangian particle, we attach energy spectra discretised in logarithmic momentum bins, which we evolved with a Fokker-Planck approach to solve:
@@
where $N(p)$ is the momentum distribution of electrons, $Q_{\mathrm{inj}}(p,t)$ is a time-dependent source term for the injection of electrons (in our case, only allowed for Diffusive Shock Acceleration) and $H(p)$ is the generalised cooling function $H(p)$:
$H(p)$ $=-\frac{2}{p}D_{\mathrm{pp}}+\sum_{i}\left|\frac{dp}{dt}\right|_{i},$ (2)
which sums up all relevant losses for the particle population (as in Paper I), as well as the gain by turbulent acceleration, via the $D_{\mathrm{pp}}$ term. For injecting and advecting Lagrangian tracers, we used the CRaTer code by Wittor et al. (2016), which allows us to generate the trajectories of $10^{5}$ tracers. These are injected at redshift $z=4$ in the highest resolution region of the simulation, with each tracer sampling a gas mass resolution of $5\times 10^{8}M_{\odot}$. The trajectory of tracers were updated using 110 snapshots recorded from the simulation, from $z=2$ to $z=0$. For each tracer we recorded the gas density, temperature, magnetic field, divergence and the curl of velocity as well as the Mach number of shocks, using cloud-in-cell interpolation. The Mach number is calculated from the temperature jump recorded by tracers, as in our previous work. To each tracer we assigned a simple initial power-law distribution of electrons in momentum space ($N(p,t=0)=N_{0}p^{-\delta}$, with $\delta=2$. The normalisation, $N_{0}$, was chosen in order for the total number of injected electrons to be $\phi=10^{-5}-10^{-7}$ of the density of thermal electrons). We solved the subsequent time-dependent diffusion-loss equation of relativistic electrons with the parallel ROGER [^1] https://github.com/FrancoVazza/JULIA/tree/master/ROGER Fokker-Planck solver, which is implemented in the Julia programming language [^2] https://julialang.org. We used $N_{b}=100$ momentum bins equally spaced in $\rm log(p)$, in the $p_{\rm min}\leq p\leq p_{\rm max}$ momentum range (with $P=\gamma m_{e}v$ and $p=P/(m_{e}c)$ is the normalised momentum of electrons). In all production runs we used $p_{\rm min}=10$ and $p_{\rm max}=10^{6}$ (hence $d\log(p)=0.05$). The details of the code are discussed in Vazza et al. (2023b).
@@
We also included the (re)-acceleration from diffusive shock acceleration (DSA), based on the Mach numbers recorded by tracers. The injection efficiency of CRe at shocks, especially in the low-Mach number regime is highly uncertain and cannot be derived from first principles (e.g. Bykov et al., 2019). Here we adopted the convenient parametrisation, given by Böss et al. (2023b), to reproduce the DSA acceleration efficiency proposed by Ryu et al. (2019), rescaled for the CR electron-to-proton ratio ($K_{\rm ep}$), which we fix by requiring an equal number density of supra-thermal cosmic-ray electrons and protons above the injection momentum, $K_{\rm ep}=[(m_{p}/m_{e})^{(1-\delta_{\rm inj})/2}]^{-1}$, as in  Pinzke et al. (2013). Any additional dependence between shock acceleration efficiency and the angle between the shock normal and the upstream magnetic field is neglected for simplicity. Our previous work has shown that in the ICM the majority of merger shocks are quasi-perpendicular and suitable for maximal electron acceleration  (e.g. Wittor et al., 2020; Banfi et al., 2020). While this Fokker Planck method, with the new implementation of the stochastic Fermi II acceleration based on Donnert et al. (2013)written and parallelised in Julia language, will be presented in more detailed forthcoming paper (Vazza et al, in prep.), we notice that the basic code has been already documented in our previous publications on this line of research (Vazza et al., 2021, 2023b, 2023a; Smolinski et al., 2023).
@@
We simulate the effect of Fermi-II re-acceleration via Adiabatic Stochastic Acceleration (ASA) model (Brunetti & Lazarian, 2016). This mechanism stems from the stochastic interaction of cosmic rays with diffusing magnetic field lines in super-Alfvenic turbulence. The acceleration rate used in the Fokker-Planck equation, $dp/dt$, depends on the solenoidal turbulent energy flux, $F_{\rm turb}=\frac{\rho\delta V^{3}_{\rm turb}}{2L}$, which is conserved in the Kolmogorov model of turbulence. As in Paper I, we use the gas vorticity to estimate the local level of solenoidal turbulence, $\delta V_{\rm turb}=|\nabla\times\@vec{v}|L$, where $L\approx 54\rm\leavevmode\nobreak\ kpc$. Now we write, $dp/dt\approx 4\leavevmode\nobreak\ D_{\rm pp}/p$, where $D_{\rm pp}\approx(54F_{\rm turb}\leavevmode\nobreak\ p^{2})/(c\leavevmode\nobreak\ \rho\leavevmode\nobreak\ V_{A})$, $\rho$ is the local gas density and $V_{A}$ is the local Alfven wave velocity, following Brunetti & Vazza (2020).
@@
We compensate for the effect of the finite numerical resolution of our simulation, which limits the small-scale dynamo amplification of magnetic fields, with a physically motivated rescaling of the magnetic fields strength associated with each tracer. In particular, we assumed that a fixed fraction $\eta_{B}$ ($=2-5\%$, see tests below) of the energy flux of turbulent kinetic energy gets dissipated into the amplification of magnetic fields, as in Beresnyak & Miniati (2016). To each tracer particle we assign the maximum value between the magnetic field given by the direct MHD simulation, and $B_{\rm turb}^{2}/8\pi\sim\eta_{B}F_{\rm turb}\tau\sim{1\over 2}\eta_{B}\rho\delta V_{\rm turb}^{2}$, where $\tau=L/\delta V_{\rm turb}$ is the turnover time. This magnetic field amplitude is used to compute synchrotron losses and the Fermi-II re-acceleration rate.
@@
Our Fokker-Planck method evolves the spectra of all tracers with a much more refined time stepping than for the advection of tracers. Typically we employ $\sim 100$ sub-cycling time steps between two time steps of the cosmological simulation, the exact number being set by the minimum between the acceleration and the loss timescales for particles with $\gamma\sim 10^{4}$, which contribute the most to the radio emission. Moreover, we used a limiter for the maximum turbulent velocity used for the Fermi-II acceleration, on the basis that in the presence of occasional complex gas velocity patterns, induced by oblique shocks and turbulence, the vorticity can become a biased proxy for solenoidal turbulence. In order not to keep a conservative value of the $D_{pp}$ re-acceleration coefficient, we set $\delta V_{\rm turb}/c_{s}\leq 0.4$ as a a maximum amplitude of the rms turbulent velocity used in the Fokker-Planck method.

### v2

@@ Abstract
Aims. We investigate whether the formation mechanisms of radio halos and mega radio halos differ, or whether they can be produced by different modalities of the same (re-)acceleration mechanism. Here we present results of a cosmological simulation of a disturbed galaxy cluster, with the aim to study the origin of mega radio halos.
Methods. We analysed the evolution of cosmic-ray electrons (CRe), subject to gains and losses using a Fokker-Planck solver. In particular, we included the effects of adiabatic stochastic acceleration (ASA) which is caused by the stochastic interaction of cosmic rays with diffusing magnetic field lines in super-Alfvenic turbulence. Moreover, we included shock acceleration and the seeding of CRe by galaxies.
@@ 1 Introduction
For at least three decades, it has been known that galaxy clusters host diffuse radio sources that are not associated with individual cluster galaxies (e.g. van Weeren et al., 2019, for a review). Such extended radio sources are associated with a perturbed dynamical state of the host cluster, and probe the existence of cosmic rays and magnetic fields in the intracluster medium (ICM) (e.g. Brunetti & Jones, 2014, for a review). The unprecedented sensitivity to large-scale emission provided by the Low Frequency Array (LOFAR, van Haarlem et al. 2013) has led to the discovery of very extended radio emission in clusters and between pairs of clusters (e.g. Govoni et al., 2019; Botteon et al., 2020; Osinga et al., 2021; Hoeft et al., 2021; Bonafede et al., 2021). Botteon et al. (2022) reported the LOFAR detection of diffuse radio emission in Abell 2255, covering almost the entire cluster volume, and containing classical radio halos and radio relics. At the same time, Cuciti et al. (2022) discovered a new class of diffuse radio sources beyond the volume of classical radio halos, dubbed mega radio halos (hereafter MegaRHs, in contrast to the ClassicalRHs introduced before). In four clusters, the ClassicalRH was found to be embedded in a much larger source that extends to scales of 2-3 Mpc, filling the volume of the cluster, at least up to $R_{500}$  [^1] Here with $R_{500}$ and elsewhere in the paper with $R_{200}$ and $R_{100}$, we mean the radius enclosing a total matter overdensity $500$, $200$, and $100$ times larger than the critical cosmic one. Furthermore, $M_{500}$, $M_{200}$, and $M_{100}$ correspondingly denote the total masses measured within the same radii. The radial profiles of the radio emission in these newly discovered MegaRHs show two distinct components: a brighter core region that corresponds to the ClassicalRH and, separated by a clear discontinuity, a more extended and shallower low-surface brightness component. Based on the small number of known MegaRHs, the transition from the ClassicalRHs to MegaRHs occurs at a radius that is about a half of $R_{\rm 500}$ from the cluster centre. The average emissivity of the MegaRHs is $\sim 20-25$ times lower than the emissivity of ClassicalRHs. For two MegaRHs, the combination of 50 and 140 MHz LOFAR observations yielded a relatively steep spectrum with a spectral index $\alpha\sim-1.6$.
@@
MegaRHs show promise to reveal the physical properties of the outskirts of galaxy clusters, in a $\sim 30$ times larger volume than what has been possible to date. They may also help understand particle acceleration and magnetic field amplification in very weakly collisional plasmas (e.g. Brunetti & Lazarian, 2011; Brunetti & Jones, 2014; Rincon et al., 2016; Sironi et al., 2023; Zhou et al., 2023; St-Onge & Kunz, 2018; Kunz et al., 2022), as well as to better explore the connection between reacceleration and electron injection by quasi perpendicular shocks in the ICM (ICM; e.g. Guo et al., 2014; Ha et al., 2021; Boula et al., 2024).
@@
Cluster-wide radio emission morphologies are compatible with a scenario where relativistic electrons are re-accelerated by second-order Fermi mechanisms in the turbulent ICM (e.g. Brunetti et al., 2001; Fujita et al., 2003; Cassano & Brunetti, 2005; Brunetti & Lazarian, 2016). In MegaRH, in particular, their steep radio spectrum makes this explanation by far the most plausible (Cuciti et al., 2022; Beduzzi et al., 2023; Nishiwaki et al., 2024).
@@
The theoretical challenge here is to maintain the relativistic electrons at energies that keep their synchrotron emission detectable at radio frequencies. The discontinuity in the surface brightness profile suggests a change either in the macrophysical (e.g. acceleration mechanisms) or in the microphysical properties of the plasma (e.g. acceleration efficiency, mean free path) when moving from the ClassicalRH to the MegaRH region. The scenario that we explore here is that both MegaRH and ClassicalRH are powered by turbulence (e.g. Brunetti et al., 2007; Brunetti & Jones, 2014), but their observable difference follows the significantly different properties of turbulence in the two regions. Previous studies, using cosmological simulations to interpret real observations (Botteon et al., 2022; Nishiwaki et al., 2024), suggest that radio emission in peripheral cluster regions require several percent of dissipation of kinetic turbulent energy for the amplification of magnetic fields and for the acceleration of relativistic electrons. This can happen if the turbulence in the MegaRH region has different properties than in the ClassicalRH region, and this work is also devoted to exploring these possible differences.
@@
In a first attempt (Beduzzi et al., 2023, hereafter Paper I), we investigated the lifecycle of relativistic electrons in the outskirts of galaxy clusters under realistic conditions. We calculated under which circumstances the observed volume of MegaRHs could be filled with radio-emitting electrons. However, in this first work we did not model the full ageing and re-acceleration process of CRe mixed with the ICM, nor did we simulate their synchrotron emission in any detail. Here we extend our previous work in order to quantitatively test the theoretical formation scenarios of MegaRH.
@@
A few other papers recently explored the formation of very extended radio emission in simulated clusters of galaxies, albeit with a more simplified approach than the one we have used for this work. By relying on the classical transit time damping (TTD) acceleration (in which the re-acceleration is powered by magneto-sonic waves excited by the compressive part of the turbulent kinetic power; e.g. Cassano & Brunetti 2005; Brunetti et al. 2007; Donnert et al. 2013), Böss et al. (2023a) simulated the emergence of diffuse radio emission in a suite of constrained simulations of the Local Universe to study the production of radio emission from turbulent re-acceleration in massive clusters, such as Coma. On the other hand, Nishiwaki et al. (2024) explored the adiabatic-stochastic-(re-)acceleration scenario (Brunetti & Lazarian, 2011; Brunetti & Vazza, 2020) and used a Fokker-Planck method to evolve the distribution of relativistic electrons in a perturbed cluster from the same suite of magnetohydrodynamics (MHD) cosmological simulations used in our Paper I. Limited to the analysis of a single cluster snapshot from the simulation, they found that if a conservative fraction ($\sim 1\%$) of solenoidal turbulent energy is dissipated for magnetic energy and electron re-acceleration, a large fraction of peripheral cluster volume can produce detectable radio emission.
@@
In this work, we apply a full Fokker-Planck solver over a population of Lagrangian tracers. This will be done for about a hundred root-grid time steps (approximately yielding a $\Delta t=100\leavevmode\nobreak\ \rm Myr$ time step) of the simulation, in order to test whether the complex cycles of (re-)acceleration experienced by relativistic electrons can reproduce observational properties of MegaRH.
@@ 2.1 Cosmological simulation: ENZO
Continuing on from Paper I, we analysed the same cosmological simulation of a cluster of galaxies, produced with ENZO code (Bryan et al., 2014), for which our previous work predicted significant chances of re-acceleration of relativistic electrons, in both the Classical and Mega RH regions. The simulation includes the self-gravity for ordinary and dark matter, radiative equilibrium cooling, but no other galaxy formation-related physics.
No feedback processes were needed as the resolution is not high enough to produce strong cooling flow, and the cluster is undergoing frequent merging to keep it from cooling castrophically. Magnetic fields are evolved using ideal MHD using the hyperbolic cleaning method (Bryan et al., 2014). At the beginning of the simulation, at a redshift of $z=40$, we injected a uniform magnetic field of $\rm B_{0}=0.4$ nG in all directions. The most massive cluster at $z=0$ has a mass of $\rm M_{100}=3.8\times 10^{14}\ M_{\odot}$ and a virial radius of $\rm R_{100}=1.5$ Mpc. The simulation has been performed on a computational grid of $128^{3}$ cells, namely $(100\rm\leavevmode\nobreak\ Mpc/h^{-1})^{3}$, with dark matter particles ($\rm m_{DM0}=3\times 10^{10}M_{\odot}$ of mass resolution). A zoomed-in region where the cluster forms (with size $(16\,{\rm Mpc})^{3}$) was further refined with four additional levels of ×$\times 2$ refinement in spatial resolution each (and ×$\times 2^{3}$ refinement in mass resolution of the dark matter component, at each refinement step) with nested regions of decreasing size. The entire cluster region is thus resolved with a maximum dark matter mass resolution of $m_{\rm DM}=7.3\times 10^{6}M_{\odot}$ (i.e. $m_{\rm DM0}/8^{4}$), and with a uniform spatial resolution of $70$ kpc/cell. On top of this, we also allowed for two extra levels of adaptive mesh refinement within the innermost region by refining on local gas or dark matter overdensities, up to a final maximum resolution of $\approx 12\rm\leavevmode\nobreak\ kpc/h=18$ kpc (comoving).
@@
The assumed cosmology is a flat $\Lambda$CDM model with $H_{0}=67.8\,\rm km/s/Mpc$, $\Omega_{M}=0.308$, $\Omega_{\Lambda}=0.692$, $\Omega_{b}=0.0478$, and $\sigma_{8}=0.811$.
@@
Figure 2: Evolving momentum spectra of all populations of relativistic electrons, for three physical models (C=only cooling; CT=cooling and turbulent re-acceleration; CST=cooling, turbulent-reacceleration and shock acceleration).
@@ 2.2 Simulations of cosmic-ray electron spectra: CRATER & ROGER
We modelled the spatial evolution of relativistic electrons in post-processing by sampling of a distribution of discrete Lagrangian particles, created by sampling a fixed gas mass resolution (approximately of $5\cdot 10^{8}M_{\odot}$) in the grid of density values produced by the ENZO simulation. Once injected, the trajectories of tracers are evolved forward in time, with a simple time integrator ${r}(t+dt)=\@vec{r}(t)+\@vec{v_{\rm gas}}dt$ (where ${r}(t)$ is the 3-dimensional position as a function of time and ${v_{\rm gas}}$ is the Eulerian gas velocity referred to the position of the tracer), after interpolating the 3d-velocity field with a cloud-in-cell spatial interpolation method, as explained in Wittor et al. (2017). To each Lagrangian particle, we attach energy spectra discretised in logarithmic momentum bins, which we evolved with a Fokker-Planck approach to solve:
@@
where $N(p)$ is the momentum distribution of electrons, $Q_{\mathrm{inj}}(p,t)$ is a time-dependent source term for the injection of electrons (in our case, only allowed for diffusive shock acceleration; DSA) and $H(p)$ is the generalised cooling function $H(p)$:
$H(p)$ $=-\frac{2}{p}D_{\mathrm{pp}}+\sum_{i}\frac{dp}{dt}_{i},$ (2)
which sums up all relevant losses for the particle population (as in Paper I: ionisation and Coulomb losses, adiabatic losses, synchrotron and Inverse Compton losses, e.g. Sarazin 1999), as well as the gain by turbulent acceleration, via the $D_{\mathrm{pp}}$ term. While we give a detailed explanation of how the $D_{\mathrm{pp}}$ turbulent re-acceleration term is computed below (as it is central to our study), we give an overview of the rest of the adopted cooling terms in Appendix A. For injecting and advecting Lagrangian tracers, we used the CRaTer code by Wittor et al. (2016), which allows us to generate the trajectories of $10^{5}$ tracers. These are injected at redshift $z=4$ in the highest resolution region of the simulation, with each tracer sampling a gas mass resolution of $5\times 10^{8}M_{\odot}$. The trajectory of tracers were updated using 110 snapshots recorded from the simulation, from $z=2$ to $z=0$. For each tracer we recorded the gas density, temperature, magnetic field, divergence and the curl of velocity as well as the Mach number of shocks, using cloud-in-cell interpolation. The Mach number is calculated from the temperature jump recorded by tracers, as in our previous work. To each tracer we assigned a simple initial power-law distribution of electrons in momentum space, that is, $N(p,t=0)=N_{0}p^{-\delta}$, with $\delta=2$. The normalisation, $N_{0}$, was chosen in order for the total number of injected electrons to be $\phi=10^{-5}-10^{-7}$ of the density of thermal electrons. This range of values is compatible with the number density of cosmic ray electrons which can be inferred based on the extent and power of observed radio halos (e.g. Vazza & Botteon, 2024, for a recent review). We solved the subsequent time-dependent diffusion-loss equation of relativistic electrons with the parallel ROGER [^2] https://github.com/FrancoVazza/JULIA/tree/master/ROGER Fokker-Planck solver, which is implemented in the Julia programming language [^3] https://julialang.org. We used $N_{b}=100$ momentum bins equally spaced in $\rm log(p)$, in the $p_{\rm min}\leq p\leq p_{\rm max}$ momentum range (with $P=\gamma m_{e}v$ and $p=P/(m_{e}c)$ is the normalised momentum of electrons). In all production runs we used $p_{\rm min}=10$ and $p_{\rm max}=10^{6}$ (hence $d\log(p)=0.05$). The details of the code are discussed in Vazza et al. (2023b).
@@
We also included the (re-)acceleration from DSA, based on the Mach numbers recorded by tracers which enters in the solved Fokker-Planck equations of Eq. 1 as an additional source term which is only introduced whenever strong enough shocks are recorded by the tracers, $Q_{\rm inj}(p,t)$.
@@
The injection efficiency of CRe at shocks, especially in the low-Mach number regime is highly uncertain and cannot be derived from first principles (e.g. Bykov et al., 2019). Moreover, more complex and realistic theoretical recipes to model the injection of electrons arising from the contribution of shock drift acceleration in low-Mach number and oblique shocks can be considered (e.g. Amano & Hoshino, 2022), and thus our prescription for $Q_{\rm inj}(p,t)$ is meant to be a first, simplistic choice to test our algorithm. Here we adopted the convenient parametrisation, given by Böss et al. (2023b), to reproduce the DSA acceleration efficiency proposed by Ryu et al. (2019), rescaled for the CR electron-to-proton ratio ($K_{\rm ep}$), which we fix by requiring an equal number density of supra-thermal CRe and protons above the injection momentum, $K_{\rm ep}=[(m_{p}/m_{e})^{(1-\delta_{\rm inj})/2}]^{-1}$, as in  Pinzke et al. (2013). Any additional dependence between shock acceleration efficiency and the angle between the shock normal and the upstream magnetic field is neglected for simplicity. Our previous work has shown that in the ICM the majority of merger shocks are quasi-perpendicular and suitable for maximal electron acceleration  (e.g. Wittor et al., 2020; Banfi et al., 2020). While this Fokker-Planck method, with the new implementation of the stochastic Fermi II acceleration based on Donnert et al. (2013) written and parallelised in Julia language, will be presented in more detailed forthcoming paper (Vazza et al, in prep.), we notice that the basic code has been already documented in our previous publications on this line of research (Vazza et al., 2021, 2023b, 2023a; Smolinski et al., 2023).
@@
We simulate the effect of Fermi-II re-acceleration via ASA model (Brunetti & Lazarian, 2016). This mechanism stems from the stochastic interaction of cosmic rays with diffusing magnetic field lines in super-Alfvenic turbulence. The acceleration rate used in the Fokker-Planck equation, $dp/dt$, depends on the solenoidal turbulent energy flux, $F_{\rm turb}=\frac{\rho\delta V^{3}_{\rm turb}}{2L}$, which is conserved in the Kolmogorov model of turbulence. As in Paper I, we use the gas vorticity to estimate the local level of solenoidal turbulence, $\delta V_{\rm turb}=|\nabla\times\@vec{v}|L$, where $L\approx 54\rm\leavevmode\nobreak\ kpc$. Now we write, $dp/dt\approx 4\leavevmode\nobreak\ D_{\rm pp}/p$, where $D_{\rm pp}\approx(54F_{\rm turb}\leavevmode\nobreak\ p^{2})/(c\leavevmode\nobreak\ \rho\leavevmode\nobreak\ V_{A})$, $\rho$ is the local gas density and $V_{A}$ is the local Alfven wave velocity, following Brunetti & Vazza (2020).
@@
We compensate for the effect of the finite numerical resolution of our simulation, which limits the small-scale dynamo amplification of magnetic fields, with a physically motivated rescaling of the magnetic fields strength associated with each tracer. In particular, we assumed that a fixed fraction $\eta_{B}$ ($=2-5\%$, see tests below) of the energy flux of turbulent kinetic energy gets dissipated for the amplification of magnetic fields, as in Beresnyak & Miniati (2016). To each tracer particle we assign the maximum value between the magnetic field given by the direct MHD simulation, and $B_{\rm turb}^{2}/8\pi\sim\eta_{B}F_{\rm turb}\tau\sim{1\over 2}\eta_{B}\rho\delta V_{\rm turb}^{2}$, where $\tau=L/\delta V_{\rm turb}$ is the turnover time. This magnetic field amplitude is used to compute synchrotron losses and the Fermi-II re-acceleration rate.
@@
Our Fokker-Planck method evolves the spectra of all tracers with a much more refined time stepping than for the advection of tracers. Typically we employ $\sim 100$ sub-cycling time steps between two time steps of the cosmological simulation, the exact number being set by the minimum between the acceleration and the loss timescales for particles with $\gamma\sim 10^{4}$, which contribute the most to the radio emission. Moreover, we used a limiter for the maximum turbulent velocity used for the Fermi-II acceleration, on the basis that in the presence of occasional complex gas velocity patterns, induced by oblique shocks and turbulence, the vorticity can become a biased proxy for solenoidal turbulence. In order not to keep a conservative value of the $D_{pp}$ re-acceleration coefficient, we set $\delta V_{\rm turb}/c_{s}\leq 0.4$ as a a maximum amplitude of the root mean square (rms) turbulent velocity used in the Fokker-Planck method.